# Body-Hand Modality Expertized Networks with Cross-attention for Fine-grained Skeleton Action Recognition

### v2

@@ III-F Implementations and Preprocessing method
To address the absence of hand skeleton data in the datasets[22, 23, 24, 25], we employ Mediapipe [26] for 2.5D hand skeleton extraction and hand-centric graph connections from RGB frames. We preprocessed the extracted data into the similar format with body skeleton. Because the number of people is up to 2 people in the datasets, we choose 2 hands(left, right) from the most active person in the sample. By the simple selective method, we can filter the noise from the mixture of noisy features. The extracted hand data undergoes a hand-centric preprocessing pipeline with modified sort and interpolation process. Lastly, the hand data is adjusted to match the body data dimensions (by augmenting with dummy keypoints). This ensures compatibility with existing body skeleton data, enabling cohesive multi-modal learning.
@@ IV EXPERIMENTS
In this section, we demonstrate the effectiveness and efficiency of our proposed method, BHaRNet, through extensive evaluations on large-scale skeleton-based benchmarks and a multi-modal integration task. We also conduct ablation studies to examine the impact of key design choices such as complementary loss, pooling attention, and expertized branches.
In this session, we denote BHaRNet-P as our dual-stream network with pooling attention module. BHaRNet-E refers to the Expertized Branch Model and BHaRNet-M denotes our multi-modal architecture.
@@ IV-A Expertimental Setup
We adopt the open-source DeGCN framework[21] as our skeleton baseline, which uses CTR-GCN[8] plus temporal shift modules(TSM). We changed two-stream architecture to single-stream for the better backbone to be used in our dual-stream network. Joint training is performed by fine-tuning pre-trained expertized models on the same dataset.
@@
IV-B Comparison with State-of-the-Art on Skeleton-Based Action Recognition
We first compare our skeleton-based model against existing methods on the NTU and N-UUCLA benchmarks. A  Table I compares our method with state-of-the-art methods in terms of accuracy, GFLOPs, and parameters. SkeleT[10] incorporates body, hand, and foot joints into a unified graph and achieves strong performance.
Our method attains accuracy close to other SOTA approaches like 3Mformer and SkeleT, but uses significantly fewer FLOPs and parameters. For instance, on NTU 60 X-Sub, Ours(BHaRNet-P) reaches 96.3% while keeping GFLOPs nearly half of InfoGCN’s. Although SkeleT yields marginally higher accuracy(about +0.7%), our model saves $\approx 60\%$ of FLOPs, demonstrating a strong efficiency–accuracy trade-off.
On top of that, a confusion matrix analysis(see Fig. 5) indicates that, for hand-specific actions like “snap fingers”, “make OK sign”, ”make victory sign”, and “cutting nails”. BHaRNet exhibits notably fewer misclassifications than the baseline. We attribute this to the hand branch’s specialized fine-motion features. Although our models remain highly competitive on NTU60/120, their performance on N-UCLA does not reach the same level as SOTA methods. Further discussion of these aspects will be provided in the Limitations section.
Figure 5: Confusion matrix on hand-oriented actions(10 classes to visualize). Our expertized branch model outperforms SkeleT by focusing on subtle finger articulations. Classes ’Okay sign’ and ’Victory sign’ mentioned in introduction are each index 70, 71 in the matrices.
@@ IV-C Multi-Modal Action Recognition
We next test BHaRNet in a multi-modal(skeleton + RGB) setting.  Table II summarizes the results on NTU 60/120 and PKU-MMD. By taking body-expert features from BHaRNet as a guide to RGB streams, our multi-modal approach effectively localizes relevant spatio-temporal cues in the video. Although we do not surpass SOTA methods(MMNet, PoseConv3d), the gains are consistent overall, and we remain notably more parameter- and FLOP-efficient than typical CNN-based multi-modal pipelines. This suggests that BHaRNet can integrate seamlessly with existing multi-modal architectures while preserving efficiency and capturing subtle hand or body interactions that might be lost in monolithic designs.
@@ IV-D Ablation Studies
TABLE III: Feature interaction on Joint NTU120 X-Sub. $L_{c}$ denotes for complemenatry loss. F.A. is shorten word of Fast Attention, and PAM is for Pooling Attention Module.
@@
Fast Attention [20] 92.7 92.7 2.72 4.33
@@ IV-D1 Joint Learning with Complementary Loss
Table III demonstrates the effect of training the body and hand streams jointly using a complementary loss. Without this constraint, each stream can partially overlap in feature usage, leading to suboptimal synergy. Once complementary loss is enabled, the total ensemble accuracy increases(e.g., 91.7% $\rightarrow$ 92.0% on NTU120 X-Sub using DeGCN backbone).
@@ IV-D2 Feature Interaction Modules
We compare diverse feature-fusion approaches in  Table III, including naive cross-attention, fast attention(Performer-based), and our proposed Pooling Attention Module. Fast attention improves accuracy over a simple cross-attention baseline, but pooling attention further boosts performance while controlling overall complexity. For instance, on NTU120 X-Sub, pooling attention attains 92.81%($\rightarrow$ 92.94% with complementary loss), outperforming the fast attention approach by +0.1–0.3%. Although the raw gain seems small, it is consistent across multiple seeds, and the additional overhead is modest. Hence, pooling attention can be seen as a balanced extension to fast attention that selectively focuses on the most relevant cross-modal features at each stage.
Pooling attention and the expertized branch architecture build upon the fast attention baseline. This synergy leads to consistent, if not dramatic, accuracy gains while retaining or even reducing model FLOPs. And also the performance enhancement in these modified architectures better supports for advantages of joint learning with complementary loss. Our main observation is that lightweight but well-structured fusion(pooling attention) can avoid overfitting or excessive overhead, supporting the design principle of cross-modal emphasis with experts.
@@
Model Variant ST-GCN DeGCN
@@
Body GCN 94.6 68.0 95.0
Body-expertized branch 94.6 71.0 95.1
TABLE VI: Feature interaction on Joint NTU120 X-Sub. “Cpl. Loss” means complementary loss.
@@
Score fusion w/ Cpl. Loss 92.0
Ensemble(expertized branches) 92.7
@@ IV-D3 Ablation Across Different Backbones
IV-D4 Expertized Branch vs. Single Body GCN in Multi-Modal and Single-Stream Settings
While previous ablations focused on complementary loss and feature interaction, we further investigate how an expertized body branch compares to a standard single body GCN in two scenarios: multi-modal fusion(Body + RGB) and a single-stream skeleton setting. Table V presents results on NTU120(X-Sub) when combining pose with RGB input, while Table VI shows performance in a purely skeleton-based(joint) setup.
In Table V, we compare a baseline body GCN model with our body-expertized branch, both integrated into an RGB pipeline. The body-expertized branch not only maintains strong pose-only performance(94.6%) but also provides more informative body weights for the RGB stream, boosting the RGB only accuracy from 68.0% to 71.0%. Consequently, the fused RGB+Pose performance also improves(95.1% vs. 95.0%). Although the gains are incremental, they consistently appear across multiple runs, indicating that the expertized branch transfers clearer pose cues to the RGB network than a standard body GCN does.
Table VI investigates how the expertized branches fare in a single-stream(joint) setting without multi-modal input. We compare a simple GCN ensemble(without cross-attention), the same ensemble with complementary loss, and our full ensemble of expertized branches. The expert-only ensemble(92.7%) outperforms both the naive score fusion(91.7%) and the ensemble with complementary loss(92.0%). This result suggests that even when cross-attention is not present, distinct body and hand expert branches can better exploit their respective modality cues, leading to higher overall accuracy. By explicitly preserving separate experts, the network avoids merging body and hand features prematurely, yielding a more effective final ensemble.
@@ V LIMITATION
Although BHaRNet achieves competitive results on large-scale datasets like NTU RGB+D, its accuracy on the smaller N-UCLA benchmark remains below certain specialized models. We find that the dataset’s limited size and relative scarcity of hand-centric classes reduce the benefits of our hand expert stream. In some cases, adding the under-trained hand branch even lowers the final ensemble accuracy compared to using the body stream alone. Future efforts could focus on improving the reliability of the hand stream, especially when hand-specific motions are sparse, to ensure that expert modeling remains advantageous despite limited data diversity.
@@ VI CONCLUSION
We presented BHaRNet, a dual-stream skeleton-based action recognition framework that targets fine-grained hand motions and broader body dynamics through carefully designed cross-attention and expertized branches. Extensive evaluations on skeleton-based and multi-modal setups reveal that our method achieves near-SOTA accuracy at relatively low computational costs. Ablation studies confirm that (1) a complementary loss helps each stream specialize without redundancy, and (2) our two distinct feature-interaction approaches (Pooling Attention and Expertized Branch) improve performance and computational cost, albeit for different design reasons.
@@ VII ACKNOWLEDGEMENT
We would like to thank the reviewers for their constructive feedback and suggestions. We also acknowledge the assistance of ChatGPT for editing support and grammar enhancements, which helped improve the clarity of the manuscript. We also extend our sincere gratitude to Hye-Yeon Kim for her valuable help and support throughout this work.

### v4

@@ III-F Implementations and Preprocessing method
To address the absence of hand skeleton data in the datasets[shahroudy2016ntu, liu2019ntu120, wang2014nucla, liu2017pku-mmd], we employ Mediapipe [lugaresi2019mediapipe] for 2.5D hand skeleton extraction and hand-centric graph connections from RGB frames. We preprocessed the extracted data into a similar format with body skeleton. Because there is a limit of 2 people in the datasets, we choose 2 hands(left, right) from the most active person in the sample. By the simple selective method, we can filter the noise from the mixture of noisy features. The extracted hand data undergoes a hand-centric preprocessing pipeline with modified sort and interpolation process. Lastly, the hand data is adjusted to match the body data dimensions (by augmenting with dummy keypoints). This ensures compatibility with existing body skeleton data, enabling cohesive multi-modal learning.
@@ IV EXPERIMENTS
In this section, we demonstrate the effectiveness and efficiency of our proposed method through extensive evaluations on large-scale skeleton-based benchmarks and a multi-modal integration task. We also conduct ablation studies to examine the impact of key design choices such as complementary loss, pooling attention, and expertized branches. We denote BHaRNet-P as our dual-stream network with pooling attention module. BHaRNet-E refers to the expertized branch model and BHaRNet-M denotes our multi-modal architecture.
@@ IV-A Expertimental Setup
We adopt the open-source DeGCN framework[myung2024degcn] as our skeleton baseline, which uses CTR-GCN[chen2021ctrgcn] plus temporal shift modules(TSM). We changed two-stream architecture to single-stream for the better backbone to be used in our dual-stream network. Joint training is performed by fine-tuning pre-trained expertized models on the same dataset.
@@
IV-B Skeleton-Based Action Recognition
We first compare our skeleton-based model against existing methods on the NTU 60/120 and N-UCLA benchmarks.  Table I compares our method with state-of-the-art methods in terms of accuracy, FLOPs, and parameters. SkeleT incorporates body, hand, and foot joints into a unified graph and achieves strong performance.
Our method attains accuracy close to other SOTA approaches like 3Mformer and SkeleT, but uses significantly fewer FLOPs and parameters. For instance, on NTU 120 X-Sub, Ours(BHaRNet-P) reaches 94.3% while keeping GFLOPs nearly half of BlockGCN’s. Although SkeleT yields marginally higher accuracy, our model saves $\approx 60\%$ of FLOPs, demonstrating a strong efficiency–accuracy trade-off.
On top of that, a confusion matrix analysis(see Figure 5) indicates that, for hand-specific actions like “OK sign” and “Victory sign”. BHaRNet-E exhibits notably fewer misclassifications than SkeleT. We attribute this to the hand branch’s specialized fine-motion features. Although our models remain highly competitive on NTU 60/120, their performance on N-UCLA does not reach the same level as SOTA methods. Further discussion of these aspects will be provided in Section V.
Figure 5: Confusion matrix on hand-oriented actions(10 classes to visualize). Ours(BHaRNet-E) outperforms SkeleT by focusing on subtle finger articulations. Classes “Okay sign” and “Victory sign” mentioned in introduction are each index 70, 71 in the matrices.
@@ IV-C Multi-Modal Action Recognition
We next test BHaRNet in a multi-modal(skeleton + RGB) setting.  Table II summarizes the results on NTU 60/120 and PKU-MMD. By taking body-expert features from BHaRNet-E as a guide to RGB streams, our multi-modal approach effectively localizes relevant spatio-temporal cues in the video. Although we do not surpass SOTA methods(MMNet, PoseConv3D), the gains are consistent overall, and we remain notably more parameter- and FLOP-efficient than typical CNN-based multi-modal pipelines. This suggests that our model(BHaRNet-M) can integrate seamlessly with existing multi-modal architectures while preserving efficiency and capturing subtle hand or body interactions that might be lost in monolithic designs.
@@ IV-D Ablation Studies
TABLE III: Feature interaction on Joint NTU120 X-Sub. $L_{cpl}$ denotes for complemenatry loss. F.A. is shorten word of Fast Attention, and PAM is for Pooling Attention Module.
@@
Fast Attention [choromanski2020performeratt] 92.7 92.7 2.72 4.33
@@ IV-D1 Joint Learning with Complementary Loss
Table III demonstrates the effect of training the body and hand streams jointly using a complementary loss. Without this constraint, each stream can partially overlap in feature usage, leading to suboptimal synergy. Once complementary loss is enabled, the total ensemble accuracy increases(e.g. 91.7% $\rightarrow$ 92.0% on NTU120 X-Sub).
@@ IV-D2 Feature Interaction Modules
We compare diverse feature-fusion approaches in  Table III, including naive cross-attention, fast attention, and our proposed Pooling Attention Module. Fast attention improves accuracy over a simple cross-attention baseline, but pooling attention further boosts performance while controlling overall complexity. For instance, on NTU120 X-Sub, pooling attention attains 92.8%($\rightarrow$ 92.9% with complementary loss), outperforming the fast attention approach. Although the raw gain seems small, it is consistent across multiple seeds, and the additional overhead is modest. Hence, pooling attention can be seen as a balanced extension to fast attention that selectively focuses on the most relevant cross-modal features at each stage.
Pooling attention and expertized branches are built on the dual-stream baseline with fast attention. This synergy leads to consistent, if not dramatic, accuracy gains while retaining or even reducing model FLOPs. And also the performance enhancement in these modified architectures better supports for advantages of joint learning with complementary loss. Our main observation is that lightweight but well-structured fusion(pooling attention) can avoid overfitting or excessive overhead, supporting the design principle of cross-modal emphasis with experts.
@@
Model Variant ST-GCN[yan2018stgcn] DeGCN
@@
Body GCN 94.3 68.0 95.0
Body-expertized branch 94.3 71.0 95.1
TABLE VI: Feature interaction within a pure Dual-stream Architecture on Joint NTU120 X-Sub.
@@
Score fusion w/ $L_{cpl}$ 92.0
BHaRNet-E(Expert-only) 92.7
@@ IV-D3 Ablation across Different Backbones
IV-D4 Expertized Branch vs. Single Body GCN
While previous ablations focused on complementary loss and feature interaction, we further investigate how an expertized body branch compares to a standard single body GCN in two scenarios: multi-stream fusion with RGB and a single-stream skeleton setting. Table V presents results on NTU120(X-Sub) when combining pose with RGB input, while Table VI shows performance in a purely skeleton-based setup.
In Table V, we compare a baseline body GCN model with our body-expertized branch, both integrated into an RGB pipeline. The body-expertized branch not only maintains strong pose-only performance(94.3%) but also provides more informative body weights for the RGB stream, boosting the RGB-only accuracy. Consequently, the fused RGB+Pose performance also improves. Although the gains are incremental, they consistently appear across multiple runs, indicating that the body-expertized branch transfers clearer pose cues to the RGB network than a standard body GCN does.
Table VI investigates how the expertized branches perform in a pure dual-stream architecture(Figure 2). We compare a simple GCN ensemble, the same ensemble from joint training with complementary loss, and our full ensemble of expertized branches. The expert-only ensemble in BHaRNet-E outperforms both the naive score fusion and the ensemble with complementary loss. This result suggests that even when cross-attention is not present, distinct body and hand expert branches can better exploit their respective modality cues, leading to higher overall accuracy. By explicitly preserving separate experts, the network avoids merging body and hand features prematurely, yielding a more effective final ensemble.
@@ V LIMITATION
Although our model achieves competitive results on large-scale datasets, its accuracy on the smaller N-UCLA benchmark remains below certain specialized models. We find that the dataset’s limited size and relative scarcity of hand-centric classes reduce the benefits of our hand expert stream. In some cases, adding the under-trained hand branch even lowers the final ensemble accuracy compared to using the body stream alone. Future efforts could focus on improving the reliability of the hand stream, especially when hand-specific motions are sparse, to ensure that expert modeling remains advantageous despite limited data diversity.
@@ VI CONCLUSION
We presented BHaRNet, a dual-stream skeleton-based action recognition framework that targets fine-grained hand motions and broader body dynamics through carefully designed cross-attention and expertized branches. Extensive evaluations on skeleton-based and multi-modal setups reveal that our method achieves near-SOTA accuracy at relatively low computational costs. Ablation studies confirm that (1) a complementary loss helps each stream specialize without redundancy, and (2) our two distinct feature-interaction approaches (Pooling Attention Module and Expertized Branches) improve performance and computational cost, albeit for different design reasons.
@@ VII ACKNOWLEDGEMENT
We would like to thank the reviewers for their constructive feedback and suggestions. We also acknowledge the assistance of ChatGPT for editing support and grammar enhancements, which helped improve the clarity of the manuscript. We also extend our sincere gratitude to Hye-Yeon Kim for her valuable help and support throughout this work. This work was supported by NST grant (CRC 21011, MSIT), IITP grant (RS-2023-00228996, RS-2024-00459749, RS-2025-25443318, RS-2025-25441313, MSIT) and KOCCA grant (RS-2024-00442308, MCST).